How do you derive the equation τ = k θ for a torsion spring? τ is torque, k is torsion spring’s constant, and θ is displacement in radians. Thanks!  


1
Expert's answer
2021-12-15T22:33:37-0500

Torsional pendulum

Simple harmonic motion equation written as


"I\frac{d^2\theta}{dt^2}=c\theta"

We know that

Torque

"\tau=I\frac{d^2\theta}{dt^2}"

Both equation equal we can written as

"\tau=c\theta"

Now c=k

"\tau=k\theta"


Where "\theta" =displacement (rad)